OpenStudy (sleepyjess):

What is 5 pi over 4 radians in degrees? 45° 135° 225° 315°

OpenStudy (anonymous):

pi radians=180 degree

OpenStudy (anonymous):

\[\frac{ 5 \pi }{ 4 }=\frac{ 5 }{ 4 }*180=?\]

OpenStudy (sleepyjess):

225

OpenStudy (anonymous):

correct.

OpenStudy (sleepyjess):

Thank you!